Estimate 247^9 without using a calculator. I asked my teacher and she said a hint was 200 < 247 < 300. I'm still really stuck. Please help?
Answer by math_tutor2020(3835)


200 < 247 < 300
2009 < 2479 < 3009

Raising each part to the 9th power will not flip the inequality signs.
To estimate 2479, we need to compute 2009 and 3009
Which in turn means we need 29 = 512 and 39 = 19683
Both of which can be computed without a calculator when given enough time.

2009 = (2*102)9
2009 = 29*(102)9
2009 = 512*102*9
2009 = 512*1018
2009 = 5.12*102*1018
2009 = 5.12*102+18
2009 = 5.12*1020
This massive number is 512 followed by 18 zeros.

3009 = (3*102)9
3009 = 39*(102)9
3009 = 19683*1018
3009 = 1.9683*104*1018
3009 = 1.9683*104+18
3009 = 1.9683*1022
This massive number is 19683 followed by 18 zeros.

We go from this
2009 < 2479 < 3009
to this
5.12*1020 < 2479 < 1.9683*1022
Since the numbers involved are so large, I recommend to keep things in scientific notation.